# The $L_{p}$-error rate for randomized quasi-Monte Carlo self-normalized importance sampling of unbounded integrands

### v1

@@ 5 Numerical experiments
5.1 A toy Bayesian inverse problem with linear forward problem
Consider a Bayesian inverse problem where the model is given by $\bm{y}=\mathcal{G}(\bm{z})+\tilde{\epsilon}$ with observed data $\bm{y}\in\mathbb{R}^{J}$, unknown parameter $\bm{z}\in\mathbb{R}^{d}$, response operator $\mathcal{G}$, and noise vector $\tilde{\epsilon}\sim\mathcal{N}(0,(1/n)I_{J})$ where $I_{J}$ is a $J$-dimensional identity matrix and $n$ is the level of noise. The likelihood is proportion to $\exp\left\{-n\Psi(\bm{z})\right\}$ with $\Psi(\bm{z})=\frac{1}{2}\|\bm{y}-\mathcal{G}(\bm{z})\|^{2}$.
Our aim is to estimate $\pi(f)=E_{\pi}[f(\bm{Z})]$, where $\pi$ denotes the posterior distribution of the unknown parameter $\bm{z}$. With a Gaussian prior $\pi_{0}=N\left(\bm{\mu_{0}},\bm{\Sigma}_{0}\right)$, the posterior density is
@@ 5.1 A toy Bayesian inverse problem with linear forward problem
which implies that the posterior expectation of interest can be expressed as
$\mathbb{E}_{\pi}[f(\bm{Z})]=\frac{\int_{\mathbb{R}^{d}}f(\bm{z})\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})d\bm{z}}{\int_{\mathbb{R}^{d}}\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})d\bm{z}}.$
Following [he2024], we consider a toy example to analyze the effect of the model parameters. Let $\mathcal{G}(\bm{z})=\bm{z}+\lambda\mathcal{F}(\bm{z})$ with $\lambda>0$ and $\mathcal{F}(\bm{z})=\left(z_{1}e^{-z_{1}^{2}},\ldots,z_{d}e^{-z_{d}^{2}}\right)^{\mathrm{T}}$. For simplicity, we take $\bm{y}=0$, which gives
@@
Note that $\Psi(\bm{z})\geq\frac{1}{2}\|\bm{z}\|^{2}$. By Example 4.9 of [he2024], $\Psi(\bm{z})$ satisfies Assumption 1 with arbitrarily small $M_{\Psi}>0$. Hence, the likelihood $l(\bm{z})$ satisfies Assumption 1 with $M_{l}=-\frac{n}{2}+M_{\Psi}<0$. For the Gaussian prior $N\left(\bm{\mu_{0}},\bm{\Sigma}_{0}\right)$, we set $\bm{\mu_{0}}=\bm{1}$ and $\bm{\Sigma}_{0}=I_{d}$. Therefore, $M_{\pi}=-\frac{1}{2}-\frac{n}{2}+M_{\Psi}<0$.
Note that $\bm{\mu}^{*}=\mathbf{0}$ is the global minimizer of $\Psi(\bm{z})$, and $\bm{\Sigma}^{*}=\left(\nabla^{2}\Psi\left(\bm{\mu}^{*}\right)\right)^{-1}=\kappa I_{d}$ with $\kappa=(1+\lambda)^{-2}$. In our experiments, we vary the model parameter $\kappa$, the dimension $d$, and the moment parameter $p$ to show the performance of RQMC-SNIS estimator. We consider five proposals as follows:
@@
  • tIS0: Linear t proposal, $\bm{\mu}=\mathbf{1}$, $\bm{\Sigma}=I_{d}$ and $v=5$;
  • tIS: Linear t proposal, $\bm{\mu}=\mathbf{0}$, $\bm{\Sigma}=(\kappa/n)I_{d}$ and $v=5$.
@@
For LapIS, we have $M_{L}=\kappa/n$ and $M_{q}=\frac{n}{2\kappa}$, then for any $\varepsilon_{1}>0$,
$M^{\mathrm{LapIS}}_{\omega f}=M_{f}+\frac{n}{2\kappa}-\frac{1}{2}-\frac{n}{2}+M_{\Psi}=\frac{n}{2\kappa}\left(1-\frac{n+1}{n}\kappa\right)+\varepsilon_{1}.$
If $\kappa>\frac{n}{n+1}$, then, $M^{\mathrm{LapIS}}_{\omega f,\tau}=M^{\mathrm{LapIS}}_{\omega f}M_{L}\leq 0$. Otherwise, $M^{\mathrm{LapIS}}_{\omega f,\tau}>0$ which is not arbitrarily small. For tIS and tIS0, $M_{q}$ is arbitrarily small, then,
@@
Hence, for any $p$ and $\kappa$, $M^{\mathrm{tIS0}}_{\omega f,\tau}=M^{\mathrm{tIS}}_{\omega f,\tau}<0$. For PriorIS and ODIS, we also have $M^{\mathrm{PriorIS}}_{\omega f,\tau}=M^{\mathrm{ODIS}}_{\omega f,\tau}<0$ for any $p$ and $\kappa$. Therefore, by Theorem 3, for LapIS proposal with $\kappa>\frac{n}{n+1}$ and other four proposals, we can get $\mathcal{O}(N^{-1+\epsilon})$ for $L_{p}$-error.
@@ 5.2 Bayesian logistic regression
We consider a Bayesian logistic regression model following [he2023]. Let $\bm{X}\in\mathbb{R}^{m\times d}$ denote the observation matrix of predictor variables with $\bm{X}_{i}\in\mathbb{R}^{1\times d}$ the $i$-th row, and let $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{d})^{T}\in\mathbb{R}^{d\times 1}$ denote the regression parameters. The logistic model is expressed as $\mathbb{P}\left(y_{i}=1\right)=1/(1+\exp(-\bm{X}_{i}\bm{\beta})),i=1,2,\ldots,m$. The log-likelihood function for $\bm{Y}=(y_{1},y_{2},\ldots,y_{m})$ can be derived as follows:
$\log l(\bm{\beta})=\log\mathbb{P}\left(\bm{Y}\mid\bm{X},\bm{\beta}\right)=\bm{\beta}^{T}\bm{X}^{T}\bm{Y}-\sum_{i=1}^{m}\log\left(1+\exp\left(\bm{\beta}^{T}\bm{X}_{i}\right)\right).$
Assume a standard Gaussian prior $\bm{\beta}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$. We consider the Pima Indian dataset with $m=392,d=9$ [windle2013]. Following [he2023], we compare two proposal distributions:
@@
    $\bm{\mu}_{\star}=\sum_{i=1}^{m}\left(Y_{i}-\frac{\exp\left(\bm{X}_{i}^{T}\bm{\mu}_{\star}\right)}{1+\exp\left(\bm{X}_{i}^{T}\bm{\mu}_{\star}\right)}\right)\bm{X}_{i},$
@@
    $\bm{\Sigma}_{\star}=\left(\bm{I}_{d}-\nabla^{2}F\left(\bm{\mu}_{\star}\right)\right)^{-1},$
    where
    $\nabla^{2}F\left(\bm{\mu}_{\star}\right)_{jk}=-\sum_{i=1}^{m}\frac{\bm{X}_{ij}\bm{X}_{ik}\exp\left(\bm{X}_{i}^{T}\bm{\mu}_{\star}\right)}{\left(1+\exp\left(\bm{X}_{i}^{T}\bm{\mu}_{\star}\right)\right)^{2}}.$
@@
where we have use the fact that $\lambda_{\max}(\bm{\Sigma}_{\star}^{-1})>1$ by Theorem 5.3 of [he2023]. However, we cannot guarantee that $\alpha^{\mathrm{LapIS}}/M^{\mathrm{LapIS}}_{\omega f,\tau}>1$, and consequently our theoretical results do not apply to LapIS in this model. On the other hand, for tIS we have $M^{\mathrm{tIS}}_{\omega f,\tau}=-\frac{1}{2}\lambda_{\max}(\bm{\Sigma}_{\star})+\varepsilon<0$ for any $\varepsilon>0$. The tIS proposal still achieves an $L_{p}$-error rate of order $\mathcal{O}(N^{-1+\epsilon})$. Figure 4 illustrates the estimate of $L_{p}$-error for $p=\{1,2,5\}$. It can be observed that LapIS performs worse than tIS, whereas the convergence rates of tIS remain consistently close to $\mathcal{O}(N^{-1})$ across different values of $p$.

### v2

@@ 5 Numerical experiments
5.1 Bayesian inverse problem
Consider a Bayesian inverse problem with the forward model $\bm{y}=\mathcal{G}(\bm{z})+\bm{\varsigma}$, where $\bm{y}\in\mathbb{R}^{d}$ is the observed data, $\bm{z}\in\mathbb{R}^{d}$ is the unknown parameter, $\mathcal{G}$ is the response operator, and $\bm{\varsigma}\in\mathbb{R}^{d}$ is the noise vector. Assume that $\bm{\varsigma}\sim\mathcal{N}(\bm{0},(1/n)I_{d})$, where $n$ is the level of noise. The likelihood $l(\bm{z})$ is proportion to $\exp\left\{-n\Psi(\bm{z})\right\}$ with $\Psi(\bm{z})=\frac{1}{2}\|\bm{y}-\mathcal{G}(\bm{z})\|^{2}$. Let $\pi_{0}(\bm{z})$ be prior distribution. By Bayes’ formula, the posterior distribution is given by
@@ 5.1 Bayesian inverse problem
where the denominator is usually an intractable constant.
Our aim is to estimate the posterior expectation of interest given by
$\pi(f):=\mathbb{E}_{\pi}[f(\bm{Z})]=\frac{\int_{\mathbb{R}^{d}}f(\bm{z})\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})d\bm{z}}{\int_{\mathbb{R}^{d}}\exp(-n\Psi(\bm{z}))\pi_{0}(\bm{z})d\bm{z}}.$
Following [he2024], we consider a Gaussian prior $\pi_{0}=N\left(\bm{\mu_{0}},\bm{\Sigma}_{0}\right)$ and a linear mapping with a small nonlinear perturbation as the response operator to analyze the effect of the model parameters. Specifically, we take $\mathcal{G}(\bm{z})=\bm{z}+\lambda\mathcal{F}(\bm{z})$ with $\lambda>0$ and $\mathcal{F}(\bm{z})=\left(z_{1}e^{-z_{1}^{2}},\ldots,z_{d}e^{-z_{d}^{2}}\right)^{\mathrm{T}}$. For simplicity, we take $\bm{y}=0$, yielding
@@
Note that $\Psi(\bm{z})\geq\frac{1}{2}\|\bm{z}\|^{2}$. By Example 4.9 of [he2024], $\Psi(\bm{z})$ satisfies Assumption 1 with arbitrarily small $M_{\Psi}>0$. Hence, the likelihood $l(\bm{z})$ satisfies Assumption 1 with $M_{l}=-\frac{n}{2}+M_{\Psi}<0$. For the Gaussian prior $N\left(\bm{\mu_{0}},\bm{\Sigma}_{0}\right)$, we set $\bm{\mu_{0}}=\bm{1}$ and $\bm{\Sigma}_{0}=I_{d}$. Therefore, the posterior $\pi(\bm{z})$ has a growth rate $M_{\pi}=-\frac{1}{2}-\frac{n}{2}+M_{\Psi}<0$.
Note that $\bm{\mu}^{*}=\mathbf{0}$ is the global minimizer of $\Psi(\bm{z})$, and $\bm{\Sigma}^{*}=\left(\nabla^{2}\Psi\left(\bm{\mu}^{*}\right)\right)^{-1}=\kappa I_{d}$ with $\kappa=(1+\lambda)^{-2}$. In our experiments, we vary the model parameter $\kappa$, the dimension $d$, and the moment parameter $p$ to show the performance of various RQMC-SNIS estimators. We consider five proposals as follows:
@@
  • tIS0: Linear t proposal, $\bm{\mu}=\mathbf{1}$, $\bm{\Sigma}=I_{d}$ and the degrees of freedom $v=5$;
  • tIS: Linear t proposal, $\bm{\mu}=\mathbf{0}$, $\bm{\Sigma}=(\kappa/n)I_{d}$ and the degrees of freedom $v=5$.
@@
For LapIS, we have $M_{L}=\kappa/n$ and $M_{q}=\frac{n}{2\kappa}$. Applying Lemma 4.1, then we have
$M^{\mathrm{LapIS}}_{\omega f}=M_{f}+\frac{n}{2\kappa}-\frac{1}{2}-\frac{n}{2}+M_{\Psi}=\frac{n}{2\kappa}\left(1-\frac{n+1}{n}\kappa\right)+\varepsilon_{1},$
where $\varepsilon_{1}=M_{f}+M_{\Psi}>0$ is arbitrarily small. If $\kappa>\frac{n}{n+1}$, then, $M^{\mathrm{LapIS}}_{\omega f,\tau}=M^{\mathrm{LapIS}}_{\omega f}M_{L}\leq 0$. Otherwise, $M^{\mathrm{LapIS}}_{\omega f,\tau}>0$ which is not arbitrarily small. For tIS and tIS0, $M_{q}>0$ is arbitrarily small. Then we have
@@
Hence, for any $p$ and $\kappa$, $M^{\mathrm{tIS0}}_{\omega f,\tau}=M^{\mathrm{tIS}}_{\omega f,\tau}<0$. For PriorIS and ODIS, we also have $M^{\mathrm{PriorIS}}_{\omega f,\tau}=M^{\mathrm{ODIS}}_{\omega f,\tau}<0$ for any $p$ and $\kappa$. Therefore, by Theorem 3, for LapIS proposal with $\kappa>\frac{n}{n+1}$ and other four proposals, we can get $\mathcal{O}(N^{-1+\epsilon})$ rate for $L_{p}$-error.
@@ 5.2 Bayesian logistic regression
We consider a Bayesian logistic regression model following [he2023]. Let $\bm{X}\in\mathbb{R}^{m\times d}$ denote the observation matrix of predictor variables with $\bm{X}_{i}\in\mathbb{R}^{1\times d}$ the $i$-th row, $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{d})^{T}\in\mathbb{R}^{d\times 1}$ denote the regression parameters, and $\bm{Y}=(Y_{1},Y_{2},\ldots,Y_{m})$ denote the observation with $Y_{i}\in\{0,1\}$. The logistic model is expressed as
$\mathbb{P}\left(Y_{i}=1\right)=1-\mathbb{P}\left(Y_{i}=0\right)=\frac{\exp(\bm{X}_{i}\bm{\beta})}{1+\exp(\bm{X}_{i}\bm{\beta})},\ i=1,2,\ldots,m.$
The log-likelihood function for $\bm{Y}$ is given by
$\log l(\bm{\beta})=\bm{\beta}^{T}\bm{X}^{T}\bm{Y}-\sum_{i=1}^{m}\log\left(1+\exp\left(\bm{X}_{i}\bm{\beta}\right)\right).$
We take a standard Gaussian prior $\bm{\beta}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$. We consider the Pima Indian dataset with $m=392,d=9$ [windle2013]. Following [he2023], we compare two proposal distributions:
@@
    $\bm{\mu}_{\star}=\sum_{i=1}^{m}\left(Y_{i}-\frac{\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)}{1+\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)}\right)\bm{X}_{i}^{T},$
@@
    $\bm{\Sigma}_{\star}=\left(I_{d}-\nabla^{2}F\left(\bm{\mu}_{\star}\right)\right)^{-1},$
    where $F(\bm{\beta})=\log l(\bm{\beta})$ and
    $\nabla^{2}F\left(\bm{\mu}_{\star}\right)_{jk}=-\sum_{i=1}^{m}\frac{\bm{X}_{ij}\bm{X}_{ik}\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)}{\left(1+\exp\left(\bm{X}_{i}\bm{\mu}_{\star}\right)\right)^{2}}.$
@@
where we have use the fact that $\lambda_{\max}(\bm{\Sigma}_{\star}^{-1})>1$ by Theorem 5.3 of [he2023]. However, we cannot guarantee that $\alpha^{\mathrm{LapIS}}/M^{\mathrm{LapIS}}_{\omega f,\tau}>1$, and consequently our theoretical results do not apply to LapIS in this model. On the other hand, for tIS we have $M^{\mathrm{tIS}}_{\omega f,\tau}=-\frac{1}{2}\lambda_{\max}(\bm{\Sigma}_{\star})+\varepsilon^{\prime}$ for any $\varepsilon^{\prime}>0$. Taking a small enough $\varepsilon^{\prime}$ such that $M^{\mathrm{tIS}}_{\omega f,\tau}<0$, the tIS proposal achieves an $L_{p}$-error rate of nearly $\mathcal{O}(N^{-1})$. Figure 4 illustrates the estimate of $L_{p}$-error for $p=\{1,2,5\}$. It can be observed that LapIS performs worse than tIS, whereas the convergence rates of tIS remain consistently close to $\mathcal{O}(N^{-1})$ across different values of $p$.